thanks for help i really need this question answer

Find the solution of the IVP

x' = 1 + t + x2 + tx2, x(0) = 0.
Answer #1

xo) = 0 x = 1+ A + x² + 4x² x = (1+x²) + + (1+00²) = (1+x)(1++) dx doce = (HA) at it x2 Integrating both sides. A2 Constan. t